The nickel in a 0.3824 g sample of an ore is precipitated with dimethylglyoxime. If the ore contains 15.4% Ni, what does the precipitate of nickel dimethylglyoxime weigh in grams? (Hint: dimethyglyoxime is a bidentate ligand.)
Answer #1

Mass of Ni in sample = (Mass of sample*% Ni in sample)/100 = (0.3824*15.4)/100 = 0.0589

Moles of Ni in sample = Mass of Ni/MW = 0.0589/58.69 = 0.001

Since the ligand is bidentate, so it will combine with 2 molecules of Ni per molecule of ligand.

So, moles of ligand that precipitate = 0.001/2 = 0.0005

Mass of ligand that is precipitated = Moles of dimethylglyoxime*MW = 0.0005*116.12 = 0.0581 g

So, mass of ppt that is formed = Mass of Ni + Mass of ligand = 0.0589+0.0581 = 0.117 g